# Massive  Index  Modulation  with  Combinatorics-Free Detection  for  Integrated  Sensing  and  Communications

### v1

@@ VI-B Decoding Complexity
In this section, the decoding complexities of the SotA methods, the proposed UVD-GaBP decoder of Sec. IV, and the enhancement modules in Section V are derived, which highlight the significantly reduced complexity, that is completely independent of the prohibitive combinatorial coefficient. Specifically, the computational complexity $\mathcal{C}$ is evaluated in terms of number of floating point operations (FLOPs), which represents the cost of a single arithmetic operation of a real-valued scalar, and the complexity order $\mathcal{O}(\,\cdot\,)$ representing the leading coefficient order of each system parameter variable.
VI-B1 Brute-force ML Decoder
In a brute-force ML detection, the likelihood metric for all possible codewords of the piloted GQSM codebook $\mathcal{X}$ is evaluated and the transmit signal with the minimum error is selected, i.e.,
$\hat{\bm{x}}_{\mathrm{ML}}=\underset{\bm{x}\in\mathcal{X}}{\mathrm{argmin}}\;\!||\bm{y}-\bm{H}\bm{x}||_{2}^{2}.$ (55)
Following equation (55), the computational complexity of a brute-force decoder can be obtained as
$\mathcal{C}_{\mathrm{ML}}={\binom{N_{T}}{P}}^{\!2}\!\cdot\!(8N_{R}N_{T}+4N_{R})$ $<\mathcal{O}\Big{(}{\binom{N_{T}}{P}}^{\!2}N_{T}N_{R}\Big{)},$ (56)
which reveals that the complexity of the ML decoder is trivially dependent on the full codebook size $\binom{N_{T}}{P}^{\!2}$.
VI-B2 IQ-decoupled Vector-valued GaBP Decoder
Next, the complexity of the IQ-decoupled vector-valued GaBP (VGaBP) scheme of [40] is analyzed, which was shown to approach the optimal performance of the ML, while achieving a quadratic order complexity reduction by exploiting the half-sized decoupled i.i.d. variables of the GQSM in the IQ domain of equation (8). Following [40], the decoding complexity of the IQ-decoupled VGaBP is given by
$\mathcal{C}_{\mathrm{IQ}}$ $\!\triangleq\tau\big{(}2N_{R}[{\binom{N_{T}}{P}}(15N_{T}^{2}+15N_{T}+4)+3N_{T}^{2}-2]+4N_{T}^{2}\big{)}$
$+2N_{R}(N_{T}^{2}+2N_{T}+1)\!+\!{\binom{N_{T}}{P}}(6N_{T}^{2}+9N_{T}+2)\!+\!5N_{T}^{2},$
$~{}~{}<\mathcal{O}\Big{(}\tau{\binom{N_{T}}{P}}N_{T}^{2}N_{R}\Big{)},$ (57)
where $\tau$ is the number of MP iterations.
VI-B3 Proposed UVD-GaBP Decoder
Following the proposed MP rules, the computational complexity of Algorithm 1 without enhancements, is obtained as
$\mathcal{C}_{\mathrm{UVD}}$ $\triangleq\tau\!\cdot 4N_{R}P[2P(6N_{T}^{2}+3N_{T}-1)+6N_{T}^{2}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\;+8N_{T}N_{R}+4N_{T}+4N_{R}+1]$
$~{}\;~{}+(8N_{T}N_{R}P+4N_{R}P+10N_{T}P+4P),$
$<\mathcal{O}\big{(}\tau P^{2}N_{T}^{2}N_{R}^{2}\big{)}.$ (58)
In contrast, the complexity of the enhanced UVD-GaBP decoder of Algorithm 2 is given by
$\mathcal{C}_{\mathrm{E\text{-}UVD}}$ $\triangleq\tau\!\cdot 4N_{R}P^{2}[2P(6N_{T}^{2}+3N_{T}+2N_{T}N_{R}-1)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+6N_{T}^{2}+8N_{T}N_{R}+4N_{T}+4N_{R}+1]$
$~{}~{}\;+P(12N_{T}N_{R}P+4N_{R}P+10N_{T}P+4P),$
$<\mathcal{O}\big{(}\tau P^{3}N_{T}^{2}N_{R}^{2}\big{)}.$ (59)
@@ VI-B3 Proposed UVD-GaBP Decoder
Figure 9: Decoding complexity of the SotA and proposed UVD-GaBP algorithm for the baseline and enhanced cases, with varying number of transmit antennas $N_{T}$ and number of transmit symbols $P$, fixed $N_{R}=64$, and $\tau=100$.
It can be seen that the intractible combinatorial coefficient of the GQSM is not inherited by the proposed UVD-GaBP decoder, and instead replaced with a significantly lower polynomial-order complexity order term on the system size parameters $P,N_{T},$ and $N_{R}$.
@@
For $P=1$, the combinatorial coefficient $\binom{N_{T}}{P}$ will reduce to $N_{T}$ and hence the ML decoder enjoys a significantly low complexity, however, for actual combinatorial IM codebooks even with $P=2$ or $P=3$, the complexity of the ML decoder and the IQ-VGaBP is shown to increase to an infeasible order with respect to $N_{T}$ and $P$, especially for mMIMO scenarios with $N_{T}\geq 16$.
@@ VII Conclusion
We proposed a new paradigm for B5G massive MIMO systems enabling high-performance ISAC via leveraging the promising IM technique. The proposed scheme is enabled by a novel low-complexity piloted IM detector based on a UVD formulation of IM schemes and the GaBP framework, which was shown to achieve a decoding complexity that is completely independent of the previously prohibitive combinatorial factor. In addition, several enhancements to improve the performance of piloted IM with the UVD-GaBP is presented, in the form of transmit symbol constellation optimization and modified MP rules based on conditional PMF-based denoising and successive IC. The effectiveness of the proposed method is verified via simulation results with the GQSM as the exemplary technique, which is shown to outperform the classic MIMO multiplexing scheme in terms of the BER performance in large mMIMO systems, in addition to showing the reduced complexity of the proposed detector by showing unprescedented GQSM detection results for system sizes up to $96\times 96$ transmit and receive antennas.

### v2

@@ VI-B Decoding Complexity
In this section, the decoding complexities of the SotA methods, the proposed UVD-GaBP decoder of Sec. IV, and the enhancement modules in Section V are derived, highlighting the significantly reduced complexity, that is completely independent of the prohibitive combinatorial coefficient. Specifically, the computational complexity $\mathcal{C}$ is evaluated in terms of the number of floating point operations (FLOPs), which represents the cost of a single arithmetic operation of a real-valued scalar, and the complexity order $\mathcal{O}(\,\cdot\,)$ representing the leading coefficient order of each system parameter variable.
$\mathcal{C}_{\mathrm{ML}}$ $\!\triangleq\!{\binom{N_{T}}{P}}^{\!2}\!\!\cdot\,\!(8N_{R}N_{T}+4N_{R}),$ (56a)
$\mathcal{C}_{\mathrm{IQ}}$ $\!\triangleq\!\tau\big{(}2N_{R}[{\binom{N_{T}}{P}}(15N_{T}^{2}+15N_{T}+4)+3N_{T}^{2}-2]+4N_{T}^{2}\big{)}+2N_{R}(N_{T}^{2}+2N_{T}+1)\!+\!{\binom{N_{T}}{P}}(6N_{T}^{2}+9N_{T}+2)\!+\!5N_{T}^{2},$ (56b)
$\mathcal{C}_{\mathrm{UVD}}$ $\!\triangleq\!\tau\!\cdot 4N_{R}P[2P(6N_{T}^{2}+3N_{T}-1)+6N_{T}^{2}+8N_{T}N_{R}+4N_{T}+4N_{R}+1]+P(8N_{T}N_{R}+4N_{R}+10N_{T}+4),$ (56c)
$\!\!\!\!\!\mathcal{C}_{\mathrm{E\text{-}UVD}}$ $\!\triangleq\!\tau\!\cdot 4N_{R}P^{2}[2P(6N_{T}^{2}\!+\!3N_{T}\!+\!2N_{T}N_{R}\!-\!1)+6N_{T}^{2}\!+\!8N_{T}N_{R}+4N_{T}+4N_{R}+1]+P^{2}(12N_{T}N_{R}+4N_{R}+10N_{T}+4).\!\!\!\!$ (56d)
@@
In light of the above, Table III summarizes the decoding complexity of four detection algorithms for piloted GQSM in both the complexity order and relative complexity in FLOPs.
Figure 9: Decoding complexity of the considered algorithms, with varying number of transmit antennas $N_{T}$ and number of transmit symbols $P$, fixed $N_{R}=64$, and $\tau=100$.
For the brute-force ML detector, the likelihood metric is evaluated for all codewords of the piloted GQSM codebook $\mathcal{X}$, and the codeword with the smallest metric is selected, i.e.,
$\hat{\bm{x}}_{\mathrm{ML}}=\underset{\bm{x}\in\mathcal{X}}{\mathrm{argmin}}\;\!||\bm{y}-\bm{H}\bm{x}||_{2}^{2},\vspace{-1ex}$ (55)
where the required computational complexity in FLOPs is given by equation (56a), which is trivially dependent on the combinatoric full codebook size.
Next, the SotA IQ-decoupled vector-valued GaBP (VGaBP) scheme of [43] is analyzed, which was shown to approach the optimal performance of the ML decoder. Following [43], the decoding complexity of the IQ-decoupled VGaBP is given by equation (56b), with $\tau$ number of MP iterations, where a quadratic order complexity reduction is attained by exploiting the half-sized decoupled i.i.d. variables of the GQSM in the IQ domain of equation (8).
Finally, the decoding complexities of the two proposed algorithms are obtained by following each step of the proposed MP rules elaborated in Sections IV and V, which are respectively given by equations (56c) and (56d).
By comparing the complexity order, it can be seen that the intractible combinatorial coefficient of the GQSM is not inherited by the proposed UVD-GaBP decoders, and instead replaced with a significantly lower polynomial-order complexity order term on the system size parameters $P,N_{T},$ and $N_{R}$.
@@
For $P=1$, the combinatorial coefficient $\binom{N_{T}}{P}$ will reduce to $N_{T}$ and hence the ML decoder enjoys a significantly low complexity, however, for actual combinatorial IM codebooks even with $P=2$ or $P=3$, the complexity of the ML decoder and the IQ-VGaBP is shown to increase to an infeasible order with respect to $N_{T}$ and $P$, especially for significantly large scales of the transmitter with $N_{T}\geq 16$.
@@
While already illustrated by the behavior for the smaller values of $P=1,2,3,4$, it should also be highlighted that the complexity advantage between the non-combinatorial and combinatorial detectors becomes more significant for increasing values of $P$, which is not plotted for the sake of visualization. Therefore, when $P$ is parametrized to yield the maximum spectral efficiency, which is $P=N_{T}/2$ for piloted IM schemes and $P>N_{T}/2$ for digitally modulated IM schemes such as GQSM [30, 42], the combinatoric-free advantage of the proposed decoders will be more protuberant.
@@ VII Conclusion
We proposed a new paradigm for feasibly enabling truly massive scales of IM schemes in expectation for the B5G, by proposing a novel detection algorithm which is shown to achieve a decoding complexity that is completely independent of the previously prohibitive combinatorial factor. The proposed scheme is based on a novel UVD formulation of piloted IM signals which is generally applicable to many types of IM schemes, on which the tailored GaBP rules are designed, aided by the novel derivation of the index activation probabilities. In addition, several enhancements to improve the performance of piloted IM with the UVD-GaBP is presented, in the form of transmit symbol constellation optimization and modified MP rules based on conditional PMF-based denoising and successive IC. The effectiveness of the proposed method is verified via simulation results with the GQSM as the exemplary technique, which is shown to outperform the classical MIMO multiplexing scheme in terms of the BER performance in very large MIMO systems, in addition to showing the reduced complexity of the proposed detector by showing unprescedented GQSM detection results for system sizes up to $96\times 96$ transmit and receive antennas.